OpenStudy (anonymous):

let g(x)=2x and h(x)=x squared +4 h of g of 1

OpenStudy (gorv):

g(x)=2x h(x)=x^2+4 hog(x)=h(g(x))=h(g(1)) g(1)=1^2+4 g(1)=5 h(g(1))=h(5)=2*5=10